The area of a rectangle is 18 in square. The length is twice the width. What is the length and width of the rectangle. Thanks
Answer by Jstrasner(112)

Hey,
So for this one we need to see what we know and figure it out from there:
First, we know that the area of the rectangle is 18 in^2. This means that the length (L) multiplied by the width (W) equals 18. In other words, L x W = 18.
Second, we know that the length (L) is twice the width (W). In other words, L = 2W.
We can use this last bit of information by substituting 2W in the area equation for L:
(2W) x W = 18 => 2W^2 = 18 => (2W^2)/2 = 18/2 => W^2 = 9 => W = 3. Now we can substitute 3 for W in the other equation:
L = 2(3) => L = 6.
Therefore,
L = 6, W = 3
I hope this helps!